# Fake News in Sheep’s Clothing: Robust Fake News Detection Against LLM-Empowered Style Attacks

### v1

@@ 6.7. Case Study
To illustrate SheepDog’s potential for interpretability in real-world scenarios, we conduct a case study based on a fake news sample from the LUN test set. The baseline RoBERTa model correctly identifies the original article as fake, but mistakenly categorizes two style-transformed adversarial articles as real news under LLM-empowered style attacks. In contrast, our SheepDog approach adeptly detects the original article as fake, as well as the two adversarial counterparts. As shown in Figure 3, the article falsely asserts the efficacy of cannabis oil in treating cancer, a claim clearly at odds with established medical knowledge. By leveraging the softmax-converted probabilities derived from the attribution-level prediction scores (refer to Eq. 5), our SheepDog model effectively explains the article’s dubious nature as “false or misleading information”. Combining content-focused veracity attributions with style-agnostic training, SheepDog offers practitioners a valuable tool for reliably discerning and interpreting news veracity.
@@ 7. Conclusion and Future Work
In this paper, we address the critical aspect of style-related robustness in fake news detection. Motivated by our empirical finding on the susceptibility of state-of-the-art text-based detectors to LLM-empowered style attacks, we introduce SheepDog, a style-agnostic fake news detector that emphasizes content veracity over style. Leveraging the strengths of task-specific LM backbones and versatile general-purpose LLMs, SheepDog adopts a multi-task learning paradigm, which integrates style-agnostic training and content-focused veracity attributions. Extensive experiments on three benchmark datasets demonstrate SheepDog’s robustness and effectiveness across various style-based adversarial settings, news reframing prompts, and representative LM backbones. Moving forward, SheepDog lays a solid foundation for developing more resilient and adaptable models in the ever-changing online landscape, and demonstrates promising potential to be further extended to multi-modal scenarios.
Appendix A Dataset Description
In Section 6, we conduct extensive experiments on three widely-used real-world benchmark datasets: the FakeNewsNet benchmark (Shu et al., 2020), which consists of the PolitiFact and GossipCop datasets, and the Labeled Unreliable News (LUN) dataset (Rashkin et al., 2017). PolitiFact and LUN contain news articles about political discourse, while GossipCop focuses on celebrity gossip.
In the case of the LUN dataset, which further classifies unreliable news into three sub-categories: satire, hoax, and propaganda, we conduct binary classification between reliable (real) and unreliable (fake) news, and ensure an equal number of unreliable news from each of these fine-grained categories. The detailed statistics for each dataset are shown as follows:
Table 8. Dataset statistics.
Dataset PolitiFact GossipCop LUN
# News Articles 450 7,916 7,500
# Real News 225 3,958 3,750
# Fake News 225 3,958 3,750
Appendix B Details of LLM Prompting
B.1. Prompt Template for Baseline LLM Detectors
@@
In section 6.1.2, we select three representative LLMs as baselines for fake news detection: GPT3.5 (OpenAI, 2022) (version name: gpt-3.5-turbo), InstructGPT (Ouyang et al., 2022) (version name: gpt-3.5-turbo-instruct), and Llama2-13B (Touvron et al., 2023) (version name: Llama-2-13b-chat-hf).
@@
For GPT-3.5 and InstructGPT, we utilize their APIs from OpenAI, and set the temperature to 00 for stable veracity predictions. For Llama2-13B, we utilize the model weights from HuggingFace Transformers (Wolf et al., 2020) version 4.31.0, and set do_sample to False for greedy decoding. For all three LLMs, we adopt the following prompt to conduct zero-shot fake news detection:
{tcolorbox}
@@
[colback=black!2!white,colframe=white!50!black,boxrule=0.5mm] Question: Does the following contain real or fake news (or information)? Answer in one word with either ‘Real’ or ‘Fake’, then explain why. [news article]
@@
B.2. Prompt Template for Obtaining SheepDog’s News Reframings
@@
Recall from Section 5.1 that we generate reliable-style reframings and unreliable-style reframings with GPT3.5 for each labeled news article. Among the four prompt templates presented below, we use the first two templates to generate reliable-style reframings, and the other two to generate unreliable-style reframings.
@@
During reframing generation, we set the temperature to $0.7$, and set the maximum number of response tokens to $512$. A detailed reframing example is presented in Table 9.
{tcolorbox}
[colback=black!2!white,colframe=white!50!black,boxrule=0.5mm] Rewrite the following article in an objective and professional tone: [news article]
{tcolorbox}
[colback=black!2!white,colframe=white!50!black,boxrule=0.5mm] Rewrite the following article in a neutral tone: [news article]
{tcolorbox}
[colback=black!2!white,colframe=white!50!black,boxrule=0.5mm] Rewrite the following article in an emotionally triggering tone: [news article]
{tcolorbox}
[colback=black!2!white,colframe=white!50!black,boxrule=0.5mm] Rewrite the following article in a sensational tone: [news article]
B.3. Prompt Template for Obtaining SheepDog’s Veracity Attributions
In Section 5.3, we elicit veracity-related information based on the news content from GPT3.5, specifically by prompting the LLM to assess the labeled articles and their corresponding reframings based on a predefined set of content-focused rationales. We utilize the following prompt template:
{tcolorbox}
[colback=black!2!white,colframe=white!50!black,boxrule=0.5mm] Article: [news article]
In the process of prompting GPT3.5 for veracity attributions, we set the temperature to 00 to yield stable attributions, and set the maximum number of response tokens to $300$. We demonstrate an example in Table 10.
@@ B.3. Prompt Template for Obtaining SheepDog’s Veracity Attributions
Table 9. An example of GPT-3.5 generated reliable-style news reframing.
PROMPT: Rewrite the following article in an objective and professional tone: BREAKING: Obamas Daughter BUSTED By The Cops For Doing This At the Democratic National Convention, Michelle Obama tried to use her daughters as shining examples of just how far America has come. But the fact is that the young Obamas are pushing the limits as any teenager doesand probably even further. According to new reports, Malia Obama was spotted at a party in Marthas Vineyard recently. It was eventually broken up by police because of noise complaints. The first daughter was immediately rushed from the party by Secret Service before the cops arrived. The West Tisbury police declined to confirm if Malia was there, citing national security. But the truth surfaced anyway. This is just one incident in a string of wild behavior on Malias part. Just a month ago the first daughter came under fire when photos surfaced of her smoking marijuana at a concert. Malia is clearly using her gap year before Harvard to party and experiment. Later, Malia and Barack were spotted strolling around a pool together having a chat. Were sure he had a few choice words for her. Clearly, a lack of respect for the law runs deep in this family.
RESPONSE: Malia Obama’s Recent Activities Raise Concerns Recently, reports have emerged regarding the behavior of Malia Obama, daughter of former President Barack Obama and Michelle Obama. At the Democratic National Convention, Michelle Obama showcased her daughters as symbols of progress in America. However, it appears that the young Obamas, like many teenagers, are testing boundaries and possibly going beyond them. According to sources, Malia Obama was seen at a gathering in Marthas Vineyard that was eventually disbanded by law enforcement due to noise complaints. The first daughter was promptly escorted away from the scene by the Secret Service before the arrival of the police. West Tisbury police have neither confirmed nor denied Malia’s presence at the event, citing national security concerns. Nevertheless, the truth regarding her attendance eventually surfaced. This incident is just one in a series of instances where Malia’s behavior has attracted attention. Not long ago, she faced criticism when photos emerged showing her allegedly smoking marijuana at a concert. It appears that Malia is utilizing her gap year before attending Harvard University as a period for socializing and experimentation. Later on, both Malia and Barack Obama were observed engaging in a conversation while walking near a pool. It can be assumed that the former president had some words of advice or guidance for his daughter. It is evident that a lack of regard for the law seems to be a recurring issue within this family. In conclusion, recent events raise concerns about the behavior of Malia Obama. These incidents, including her presence at a party disbanded by the police and alleged marijuana use, suggest a period of exploration and a certain disregard for legal boundaries.
@@
Table 10. An example of content-focused veracity attributions extracted from GPT-3.5.

### v2

@@ 6.7. Case Study
To illustrate SheepDog’s potential for offering explanatory outputs during model inference via its attribution predictions (detailed in Section 5.3), we present a case study based on a fake news article from the LUN test set. As shown in Figure 3, the article falsely claims the effectiveness of cannabis oil in treating cancer, aiming to mislead readers, despite contradicting established medical knowledge. While the baseline RoBERTa detector correctly flags the original article as fake news, it misclassifies two style-transformed adversarial articles as real news. In contrast, SheepDog accurately identifies the original article as fake news, a prediction that remains consistent for its two adversarial counterparts. Remarkably, leveraging the softmax-converted probabilities from attribution-level prediction scores (Eq. 5), SheepDog consistently identifies ”false and misleading information” as the top-predicted attribution for debunking fake news. This style robustness is invaluable for practitioners seeking to comprehend the rationale behind each flagged fake news, aiding in human verification and assessment of prediction reliability.
6.8. Discussion: Why is SheepDog Yielding Greater Performance Gains on LUN?
SheepDog shows significantly greater performance improvements on LUN compared to PolitiFact and GossipCop in both adversarial (Section 6.2) and original unperturbed settings (Section 6.3). Specifically, it achieves [A] significant improvements on the original unperturbed LUN test set while maintaining performance comparable to the best baseline on PolitiFact and GossipCop (Table 5); and [B] notably greater improvements on LUN compared to PolitiFact and GossipCop on style-based adversarial test sets (Table 3).
These phenomena can be attributed to the unique style-related features of LUN, which include distinct writing styles of (1) individual news publishers and and (2) different news types. (e.g., hoax). Unlike PolitiFact and GossipCop, LUN’s publishers (i.e., news sites) in the training and test sets do not overlap (Rashkin et al., 2017), creating an inherent distribution shift between training and test data. Furthermore, as described in Section 6.1.1, the fake news articles in LUN encompass satire, hoax, and propaganda with distinctive writing styles. As a result, fake news detectors trained on LUN are expected to be more reliant on writing style.
Recall that our Observation 1 reveals the heavy reliance of existing text-based detectors on styles rather than news content for veracity prediction. Trained on LUN, the baseline models become overly reliant on styles specific to both publishers and news types, which potentially explains [A] and [B]:
  • On the original unperturbed LUN test set (Table 5), publisher-specific style features used by baseline models fail to generalize to test articles, as test articles are produced by news sites not included in the training data. In contrast, SheepDog, being style-agnostic, remains unaffected by changes in news publisher styles and yields significant improvements.
  • On the adversarial LUN test sets (Table 3), both publisher-specific and news type-specific style features utilized by baselines fail to generalize. Notably, LLM-empowered style attacks reverse the styles of both reliable and unreliable news (Section 4.1). These style variations make type-specific style features detrimental for veracity prediction. In contrast, SheepDog achieves style robustness through LLM-empowered news reframings and content-focused veracity attributions, thereby reliably detecting fake news.
@@ 7. Conclusion and Future Work
In this paper, we address the critical aspect of style-related robustness in fake news detection. Motivated by our empirical finding on the susceptibility of state-of-the-art text-based detectors to LLM-empowered style attacks, we introduce SheepDog, a style-agnostic fake news detector that emphasizes content veracity over style. Jointly leveraging the strengths of task-specific LM backbones and versatile general-purpose LLMs, SheepDog adopts a multi-task learning paradigm, which integrates style-agnostic training and content-focused veracity attribution prediction. Extensive experiments on three real-world benchmarks demonstrate SheepDog’s robustness and effectiveness across various style-based adversarial settings, news reframing prompts, and representative backbones. Moving forward, SheepDog lays a solid foundation for developing more resilient and adaptable models in the ever-changing online landscape, and demonstrates promising potential to be further extended to multi-modal scenarios.
8. Acknowledgements
This work was supported in part by the National Research Foundation Singapore, NCS Pte. Ltd. and National University of Singapore under the NUS-NCS Joint Laboratory (Grant A-0008542-00-00).
Appendix A Discussion: Effect of Reframings on LLM Style Robustness
This section investigates the effects of incorporating our style-diverse news reframings (Section 5.1) on LLM style robustness. Following the experimental setup for evaluating style robustness (Section 6), we report results on adversarial set A (formulation described in Table 4) unless otherwise specified. We explore two methods for adapting LLMs to diverse news styles: (1) in-context learning and (2) fine-tuning, with prompt templates and LLM configurations detailed in Appendix B.1. To mitigate randomness, we present averaged metrics from three runs of each approach.
In-Context Learning. Table 10 compares a zero-shot GPT-3.5 detector with the following three in-context learning variants: (1) GPT-3.5+ICL-2, which incorporates 2 randomly selected in-context samples (1 real and 1 fake). (2) GPT-3.5+ICL-2-R, which incorporates the same 2 in-context samples as GPT-3.5+ICL-2, enriched with style-diverse reframings. For each in-context sample, we randomly select one reliable-style reframing and one unreliable-style reframing. (3) GPT-3.5+ICL-4, which incorporates 4 randomly selected in-context samples (2 real and 2 fake).
Table 10. GPT-3.5 in-context learning performance.
Method PolitiFact GossipCop LUN
GPT-3.5 42.13 39.59 59.63
  + ICL-2 38.57 38.88 57.51
  + ICL-2-R 41.76 41.83 60.29
  + ICL-4 42.35 41.47 60.25
Results in Table 10 indicate three key findings: (1) LLMs benefit from more in-context samples, as GPT-3.5+ICL-4 outperforms GPT-3.5+ICL-2. (2) Incorporation of news reframings effectively enhances LLM style robustness, as shown by the consistent gains of GPT-3.5+ICL-2-R over GPT-3.5+ICL-2. (3) On the fake news detection task, LLM in-context learning does not offer clear benefits over LLM zero-shot. The long sequence lengths of news articles limit the number of in-context demonstrations, resulting in insufficient context and overly lengthy prompts that LLMs struggle to process.
Fine-Tuning. Table 11 compares a zero-shot LLaMA2-13B detector with the following two fine-tuned variants: (1) LLaMA2-13B+FT, which fine-tunes LLaMA2 using the original training articles. (2) LLaMA2-13B+FT-R, which fine-tunes LLaMA2 using the original training articles and their style-diverse reframings. For each article, we randomly incorporate one reliable-style reframing and one unreliable-style reframing as fine-tuning data.
Table 11. LLaMA2-13B fine-tuning performance.
Method PolitiFact GossipCop LUN
LLaMA2-13B 33.24 25.79 32.64
+ FT 36.96 47.83 34.31
+ FT-R 44.44 56.06 42.22
Both LLaMA2 variants are fine-tuned using QLoRA (Dettmers et al., 2023) for 1 epoch with a batch size of $16$ and a learning rate of $1\times 10^{-4}$. From Table 11, we observe: (1) Fine-tuning generally enhances task-specific LLM capabilities on fake news detection, as indicated by consistent improvements of LLaMA2-13B+FT over LLaMA2-13B. (2) Incorporating style-diverse news reframings as fine-tuning samples further improves style robustness, as indicated by consistent improvements of LLaMA2-13B+FT-R over LLaMA2-13B+FT. (3) LLaMA2-13B+FT-R still falls short compared to GPT-3.5 on the LUN adversarial test set and performs worse than task-specific detectors and fully fine-tuned LMs in Table 3. This suggests the need for larger, task-specific corpora for fully adapting LLMs to fake news detection, further highlighting our motivation for leveraging LLM general-purpose capabilities in a zero-shot manner, which are reasonably strong and more readily available.
Appendix B LLM Prompting Configurations
B.1. LLM Baselines
@@
In section 6.1.2, we select three representative LLMs as fake news detection baselines: GPT-3.5 (OpenAI, 2022) (model name: gpt-3.5-turbo-0301), InstructGPT (Ouyang et al., 2022) (model name: gpt-3.5-turbo-instruct), and LLaMA2-13B (Touvron et al., 2023) (model name: Llama-2-13b-chat-hf).
@@
For GPT-3.5 and InstructGPT, we use their APIs from OpenAI, and set the temperature to 00 for stable veracity predictions. For LLaMA2-13B, we employ model weights from HuggingFace Transformers (Wolf et al., 2020) version 4.31.0, with do_sample set to False for greedy decoding. All models adopt the following prompt for zero-shot fake news detection:
@@
Question: Does the following contain real or fake news (or information)? Answer in one word with either ‘Real’ or ‘Fake’, then explain why. [news article]
@@
B.2. Obtaining SheepDog’s News Reframings
@@
Recall from Section 5.1 that we generate reliable-style reframings and unreliable-style reframings with GPT-3.5 for each labeled news article. Among the four prompt templates presented below, we use the first two templates for reliable-style reframings, and the other two for unreliable-style reframings.
@@
During reframing generation, we set the temperature to $0.7$ and limit the maximum number of response tokens to $512$. See Table 13 for a detailed reframing example.
Rewrite the following article in an objective and professional tone: [news article]
Rewrite the following article in a neutral tone: [news article]
Rewrite the following article in an emotionally triggering tone: [news article]
Rewrite the following article in a sensational tone: [news article]
B.3. Obtaining SheepDog’s Veracity Attributions
In Section 5.3, we use an LLM to elicit auxiliary content-centric information directly related to news veracity. Specifically, we prompt the LLM to provide explanatory outputs for each fake news article in the training set, based on predefined content-focused rationales aimed at debunking fake news.
We draw from existing literature, which commonly emphasizes two key aspects: (1) sources of information (Jung et al., 2020; Zhang et al., 2018), and (2) correctness of news content (Allcott and Gentzkow, 2017; Shu et al., 2017). Inspired by these insights, we devise four rationales:
  • Source-related rationales: [A] lack of credible sources; and [B] inconsistencies with reputable sources. These rationales are informed by prior research emphasizing the importance of reputable sources in validating conclusions (Zhang et al., 2018) and correcting false information (Jung et al., 2020).
  • Content-related rationales: [C] false or misleading information; and [D] biased opinion. Fake news, defined as intentionally and verifiably false information (Allcott and Gentzkow, 2017; Shu et al., 2017), often manipulates biased opinions to exploit cognitive vulnerabilities in news consumers. Consequently, we devise [C] in alignment with the definition of fake news, and [D] to represent a common strategy employed by fake news producers.
Our four rationales cover the key aspects of source- and content-related indicators in debunking fake news. We utilize the following prompt to obtain content-focused veracity attributions, and set the temperature to 00 to ensure stability. An example is presented in Table 14.
Article: [fake news article]
Question: Which of the following problems does this article have? Lack of credible sources, False or misleading information, Biased opinion, Inconsistencies with reputable sources. If multiple options apply, provide a comma-separated list ordered from most to least related. Answer “No problems” if none of the options apply.
Appendix C Analysis on Content Consistency of LLM-Empowered News Reframings
LLM-empowered news reframings play a key role in achieving style robustness. Table 13 provides a comparison between a news article and its reliable-style reframing, illustrating concretely the impressive capability of LLMs to introduce different tones while preserving the news content.
To offer a more comprehensive assessment of content consistency between the original news article and its reframings, we investigate claim entailment for a straightforward and quantifiable estimation. Specifically, the original article should ideally entail the central factual claims within its reframing, and vice versa.
Given that news articles exhibit significantly longer sequences and more complex logical structures compared to the sentence pairs in natural language inference (NLI) benchmarks (e.g., SNLI (Bowman et al., 2015) and MultiNLI (Williams et al., 2018)), we opt against utilizing NLI models pre-trained on these benchmarks. Instead, we instruct a GPT-3.5 model to extract claims and infer claim entailment. The prompt template for claim extraction is as follows:
Extract and summarize the central factual claim in the following article. Article: [news article A]. Claim:
Querying the LLM yields a succinct summarization of the article’s central factual claim, typically composed of several sub-claims. An example response is provided as follows:
Response Example for Claim Extraction: Financial experts are concerned about the negative impact of China’s undervalued yuan on both Asia and the United States. They are calling on regional governments and the Group of 20 leaders to take action to prevent potential currency and trade wars. The experts emphasize the need for neighboring countries to urge China to relax its exchange rate controls in order to address the global current account imbalance. They also highlight the adverse effects of US monetary easing and China’s low exchange rate on emerging market economies. The experts are urging the G20 to address the currency problem at its upcoming summit in South Korea and to oppose unilateral devaluation moves and support currency stability.
With the extracted claim, we predict claim entailment using the following prompt:
Question: Does the following article entail the claim: [claim extracted from news article A]? Answer in one word with either ‘Yes’ or ‘No’. Article: [news article B].
While precise entailment assessment requires detailed analysis of each article’s nuances and specific claim wordings, evaluating the proportion of article pairs with predicted claim entailment provides a general estimation of content consistency. To ensure comprehensive coverage, we analyze two larger-scale benchmark datasets: GossipCop and LUN (dataset statistics in Table 2).
Table 12. Claim entailment (%) between original news articles and their objective-style reframings.
Dataset original-entail-objective objective-entail-original
GossipCop 86.20 89.17
LUN 89.22 87.53
As shown in Table 12, the proportions of claim entailment range from 86.20% to 89.22% across both datasets, suggesting a reasonably high consistency between the central factual claims presented in the original news articles and their reframings. This consistency further validates the effectiveness of our reframing approach in preserving the core content of news while injecting stylistic variations.
@@ Appendix C Analysis on Content Consistency of LLM-Empowered News Reframings
Table 13. An example of GPT3.5-generated reliable-style news reframing, truncated into sentences for a clearer comparison.
PROMPT: Rewrite the following article in an objective and professional tone: RESPONSE:
BREAKING: Obamas Daughter BUSTED By The Cops For Doing This Malia Obama’s Recent Activities Raise Concerns
At the Democratic National Convention, Michelle Obama tried to use her daughters as shining examples of just how far America has come. Recently, reports have emerged regarding the behavior of Malia Obama, daughter of former President Barack Obama and Michelle Obama. At the Democratic National Convention, Michelle Obama showcased her daughters as symbols of progress in America.
But the fact is that the young Obamas are pushing the limits as any teenager does and probably even further. However, it appears that the young Obamas, like many teenagers, are testing boundaries and possibly going beyond them.
According to new reports, Malia Obama was spotted at a party in Marthas Vineyard recently. It was eventually broken up by police because of noise complaints. According to sources, Malia Obama was seen at a gathering in Marthas Vineyard that was eventually disbanded by law enforcement due to noise complaints.
The first daughter was immediately rushed from the party by Secret Service before the cops arrived. The first daughter was promptly escorted away from the scene by the Secret Service before the arrival of the police.
The West Tisbury police declined to confirm if Malia was there, citing national security. West Tisbury police have neither confirmed nor denied Malia’s presence at the event, citing national security concerns.
But the truth surfaced anyway. Nevertheless, the truth regarding her attendance eventually surfaced.
This is just one incident in a string of wild behavior on Malias part. This incident is just one in a series of instances where Malia’s behavior has attracted attention.
Just a month ago the first daughter came under fire when photos surfaced of her smoking marijuana at a concert. Not long ago, she faced criticism when photos emerged showing her allegedly smoking marijuana at a concert.
Malia is clearly using her gap year before Harvard to party and experiment. It appears that Malia is utilizing her gap year before attending Harvard University as a period for socializing and experimentation.
Later, Malia and Barack were spotted strolling around a pool together having a chat. Later on, both Malia and Barack Obama were observed engaging in a conversation while walking near a pool.
Were sure he had a few choice words for her. It can be assumed that the former president had some words of advice or guidance for his daughter.
Clearly, a lack of respect for the law runs deep in this family. It is evident that a lack of regard for the law seems to be a recurring issue within this family. In conclusion, recent events raise concerns about the behavior of Malia Obama. These incidents, including her presence at a party disbanded by the police and alleged marijuana use, suggest a period of exploration and a certain disregard for legal boundaries.
@@
Table 14. An example of content-focused veracity attributions extracted from GPT3.5, given a fake news article as input.